Joanne, a recent retiree, requires $5,000 per year in extra income. She has $60,000 to invest and can invest in bonds paying 15 percent per year or in a Certificate of Deposit paying 7 percent per year. How much money should be invested in each realize exactly $5,000 in interest per year
Answer by lwsshak3(11628)

Joanne, a recent retiree, requires $5,000 per year in extra income. She has $60,000 to invest and can invest in bonds paying 15 percent per year or in a Certificate of Deposit paying 7 percent per year. How much money should be invested in each realize exactly $5,000 in interest per year
***
let x=amount invested in 15% bonds
60000-x=amount invested in 7% CD's
.15x+.07(60000-x)=5000
.15x+4200-.07x=5000
.08x=800
x=10000
60000-x=50000
amount invested in 15% bonds=$10,000
amount invested in 7% CD's=$50,000